Consecutive Integers. Find three consecutive integers such that the sum of their squares is 77.
Answer by checkley77(12844)

X^2+(X+1)^2+(X+2)^2=77
X^2+X^2+2X+1+X^2+4X+4=77
3X^2+6X+5=77
3X^2+6X+5-77=0
3X^2+6X-72=0
3(X^2+2X-24)=0
3(X+6)(X-4)=0
X-4=0
X=4, 5 & 6 ANSWERS.
PROOF:
462+5^2=662=77
16+25=36=77
77=77